# Improving Retrieval-Augmented Generation in Medicine with Iterative Follow-up Questions

### v2

@@ Abstract
The emergent abilities of large language models (LLMs) have demonstrated great potential in solving medical questions. They can possess considerable medical knowledge, but may still hallucinate and are inflexible in the knowledge updates. While Retrieval-Augmented Generation (RAG) has been proposed to enhance the medical question-answering capabilities of LLMs with external knowledge bases, it may still fail in complex cases where multiple rounds of information-seeking are required. To address such an issue, we propose iterative RAG for medicine (i-MedRAG), where LLMs can iteratively ask follow-up queries based on previous information-seeking attempts. In each iteration of i-MedRAG, the follow-up queries will be answered by a conventional RAG system and they will be further used to guide the query generation in the next iteration. Our experiments show the improved performance of various LLMs brought by i-MedRAG compared with conventional RAG on complex questions from clinical vignettes in the United States Medical Licensing Examination (USMLE), as well as various knowledge tests in the Massive Multitask Language Understanding (MMLU) dataset. Notably, our zero-shot i-MedRAG outperforms all existing prompt engineering and fine-tuning methods on GPT-3.5, achieving an accuracy of 69.68% on the MedQA dataset. In addition, we characterize the scaling properties of i-MedRAG with different iterations of follow-up queries and different numbers of queries per iteration. Our case studies show that i-MedRAG can flexibly ask follow-up queries to form reasoning chains, providing an in-depth analysis of medical questions. To the best of our knowledge, this is the first-of-its-kind study on incorporating follow-up queries into medical RAG.
@@
Preprint of an article submitted for consideration in Pacific Symposium on Biocomputing © 2024 World Scientific Publishing Co., Singapore
@@ 1 Introduction
In this work, we propose i-MedRAG, a simple and effective framework for incorporating follow-up queries into RAG. Specifically, we prompt LLMs to iteratively generate follow-up queries to search for additional information from external medical corpora. The queries and the corresponding answers generated with RAG will be used to augment the answer generation of the original question. Empirical results demonstrate the effectiveness of i-MedRAG on both open- and close-source LLMs, which show improved performance on the United States Medical Licensing Examination (USMLE) subset of MedQA and medical questions from the Massive Multitask Language Understanding (MMLU) dataset. Notably, our zero-shot i-MedRAG outperforms all previous prompt engineering and fine-tuning methods, setting a state-of-the-art accuracy for GPT-3.5 on the MedQA dataset (Figure 1). Our further analysis of the number of iterations and number of queries per iteration used in i-MedRAG reflects how its performance scales with different settings. Additionally, we present several case studies of i-MedRAG, showing how it overcomes the limitations in conventional RAG to find the correct answers.
@@ 2.1 Retrieval-Augmented Generation for Medicine
Retrieval-augmented generation (RAG) has been widely adopted in medicine. Here, we discuss several representative approaches. Almanac [21] is a system that augments LLMs with curated resources for medical guidelines and treatment recommendations, which shows improvements over the standard LLMs in six manually assessed metrics. Similarly, Low et al. [27] demonstrate the improvements of RAG-based systems for real-world clinical queries with manual evaluation. Clinfo.ai [22] is an open-source web application that answers clinical questions based on retrieved scientific literature from PubMed articles. Xiong et al. [23] conduct a benchmarking study with the MedRAG toolkit, and show the benefits of RAG in several medical multi-choice question answering datasets. There are also various biomedical literature search products [28] that use RAG to summarize the retrieved articles [29], such as OpenEvidence [^1] https://www.openevidence.com/ and ChatRWD [^2] https://www.atroposhealth.com/chatrwd. However, most of the RAG studies in medicine use the conventional architecture with only one round of retrieval. To the best of our knowledge, our study presents the first approach and evaluations on incorporating follow-up queries in RAG for medicine.
@@ 4.2 Main results
Table 4.1 shows the comparison results of i-MedRAG and other baseline approaches on MedQA using GPT-3.5. Official scores reported in previous research are used for a fair comparison. While methods with few-shot learning or model fine-tuning tend to perform better than the ones in a zero-shot setting, our i-MedRAG set a state-of-the-art performance of GPT-3.5 on MedQA without any training samples or parameter tuning. Among zero-shot approaches, i-MedRAG (69.68%) has a +4.61% performance improvement compared to the previous best record achieved by MedRAG (66.61%).
The results of generalizing i-MedRAG to more LLMs and data are presented in Table 4.2. We compare i-MedRAG with our implemented CoT and MedRAG to see if i-MedRAG can bring a consistent improvement of LLM performance in medical question answering. Similar to the results on GPT-3.5, the open-source Llama-3.1-8B also shows improved performance on MedQA with the help of i-MedRAG. While Llama-3.1-8B shows a close performance to GPT-3.5 in CoT and MedRAG settings, its performance is significantly improved with i-MedRAG, achieving an accuracy of 75.02%. The improved performance of GPT-3.5 and Llama-3.1-8B on MMLU-Med also demonstrates the generalizability of i-MedRAG to more medical data. As medical questions in MMLU-Med are less complex than the USMLE questions in MedQA, follow-up queries may not be necessary to find relevant information for the given question. Thus, it can be observed that the improvement by i-MedRAG compared to MedRAG is less significant in MMLU-Med than in MedQA.
@@
Performance of i-MedRAG on different LLMs and datasets. “Acc.” denotes the accuracy. “$\Delta$” shows the relative performance improvement compared with CoT. Model Method MedQA-USMLE MMLU-Med Average Acc. $\Delta$ Acc. $\Delta$ Acc. $\Delta$ GPT-3.5-Turbo CoT 65.04 +0.00% 72.91 +0.00% 68.98 +0.00% GPT-3.5-Turbo MedRAG 66.61 +2.41% 75.48 +3.52% 71.05 +3.00% GPT-3.5-Turbo i-MedRAG 69.68 +7.13% 76.40 +4.79% 73.04 +5.89% Llama-3.1-8B CoT 64.73 +0.00% 77.23 +0.00% 70.98 +0.00% Llama-3.1-8B MedRAG 66.54 +2.80% 78.05 +1.06% 72.30 +1.86% Llama-3.1-8B i-MedRAG 75.02 +15.90% 79.61 +3.08% 77.32 +8.93%
@@ 4.3 Scaling with iterations and queries
Figure 3: The performance of LLMs using i-MedRAG on MedQA and MMLU-Med with different hyperparameter values.
@@ 4.4 Case studies
To illustrate how i-MedRAG helps LLMs find the correct answer, we perform a case study on MedQA to show how GPT-3.5 iteratively generates follow-up queries and finds the correct answer for the original medical question. Table 4.4 shows the predictions of GPT-3.5 on a test question from MedQA with different prompt engineering approaches, with the useful evidence highlighted in yellow and the incorrect rationales marked with red highlights. The question asks about the mechanism of the drug for transitional cell carcinoma of the bladder which causes hearing loss. To solve the problem, it is important to find the exact drug based on the given context and then figure out how it causes the mentioned symptoms.
However, the CoT result shows that GPT-3.5 does not inherently contain sufficient medical knowledge to solve this problem. Instead of inferring the described drug, GPT-3.5 with CoT directly hallucinates a wrong option as the answer. While free radicals are relevant to hearing loss, their connection to the disease of the patient is unclear and not discussed. Compared to CoT which solely relies on the internal knowledge of LLMs, MedRAG provides an opportunity for LLMs to augment their answer generation with external medical knowledge. Nevertheless, the model output shows that the MedRAG system fails to retrieve useful information about the drug from medical corpora. Given the complex problem description, it is difficult for text retrievers to find the asked mechanism without knowing the drug.
@@ 5.1 Limitations
Another limitation of this work is the selection of hyperparameter values for optimal performance. As shown in Figure 3, different LLMs can have different hyperparameter settings for their optimal performance. Even for the same LLM, its optimal hyperparameters can vary based on the medical questions being processed. Thus, it is non-trivial to find the optimal hyperparameters of i-MedRAG for a new medical task, which may be inefficient for real-world deployments.
@@ 5.2 Future work
Given the limitations of i-MedRAG, we consider several potential future directions that could further improve the performance of retrieval-augmented generation for medicine. The first direction is the automation of hyperparameter selection in i-MedRAG. To reduce the laborious process of hyperparameter selection, one may use an LLM agent to dynamically determine how many follow-up queries should be asked each iteration. This can improve the efficiency and flexibility of the hyperparameter selection process. Another future direction is to improve the performance of i-MedRAG with few-shot demonstrations. While few-shot CoT prompting is demonstrated to perform better than the zero-shot counterpart[39], it is not easy to adapt such strategies to i-MedRAG as the reasoning process can be dynamically affected by the use of external corpora and retrievers. Investigating how i-MedRAG can benefit from one or few-shot samples could be a potential direction to further enhance its performance on medical question answering.

### v3

@@ Abstract
The emergent abilities of large language models (LLMs) have demonstrated great potential in solving medical questions. They can possess considerable medical knowledge, but may still hallucinate and are inflexible in the knowledge updates. While Retrieval-Augmented Generation (RAG) has been proposed to enhance the medical question-answering capabilities of LLMs with external knowledge bases, it may still fail in complex cases where multiple rounds of information-seeking are required. To address such an issue, we propose iterative RAG for medicine (i-MedRAG), where LLMs can iteratively ask follow-up queries based on previous information-seeking attempts. In each iteration of i-MedRAG, the follow-up queries will be answered by a conventional RAG system and they will be further used to guide the query generation in the next iteration. Our experiments show the improved performance of various LLMs brought by i-MedRAG compared with conventional RAG on complex questions from clinical vignettes in the United States Medical Licensing Examination (USMLE), as well as various knowledge tests in the Massive Multitask Language Understanding (MMLU) dataset. Notably, our zero-shot i-MedRAG outperforms all existing prompt engineering and fine-tuning methods on GPT-3.5, achieving an accuracy of 69.68% on the MedQA dataset. In addition, we characterize the scaling properties of i-MedRAG with different iterations of follow-up queries and different numbers of queries per iteration. Our case studies show that i-MedRAG can flexibly ask follow-up queries to form reasoning chains, providing an in-depth analysis of medical questions. To the best of our knowledge, this is the first-of-its-kind study on incorporating follow-up queries into medical RAG. The implementation of i-MedRAG is available at https://github.com/Teddy-XiongGZ/MedRAG.
@@
© 2024 The Authors. Open Access chapter published by World Scientific Publishing Company and distributed under the terms of the Creative Commons Attribution Non-Commercial (CC BY-NC) 4.0 License.
@@ 1 Introduction
In this work, we propose i-MedRAG, a simple and effective framework for incorporating follow-up queries into RAG. Specifically, we prompt LLMs to iteratively generate follow-up queries to search for additional information from external medical corpora. The queries and the corresponding answers generated with RAG will be used to augment the answer generation of the original question. Empirical results demonstrate the effectiveness of i-MedRAG on both open- and close-source LLMs, which show improved performance on the United States Medical Licensing Examination (USMLE) subset of MedQA and medical questions from the Massive Multitask Language Understanding (MMLU) dataset. Our further analysis of the number of iterations and number of queries per iteration used in i-MedRAG reflects how its performance scales with different settings. Additionally, we present several case studies of i-MedRAG, showing how it overcomes the limitations in conventional RAG to find the correct answers.
@@ 2.1 Retrieval-Augmented Generation for Medicine
Retrieval-augmented generation (RAG) has been widely adopted in medicine. Here, we discuss several representative approaches. Almanac [21] is a system that augments LLMs with curated resources for medical guidelines and treatment recommendations, which shows improvements over the standard LLMs in six manually assessed metrics. Similarly, Low et al. [27] demonstrate the improvements of RAG-based systems for real-world clinical queries with manual evaluation. Clinfo.ai [22] is an open-source web application that answers clinical questions based on retrieved scientific literature from PubMed articles. Xiong et al. [23] conduct a benchmarking study with the MedRAG toolkit, and show the benefits of RAG in several medical multi-choice question answering datasets. There are also various biomedical literature search products [28] that use RAG to summarize the retrieved articles [29], such as OpenEvidence [^1] https://www.openevidence.com/ and ChatRWD [^2] https://www.atroposhealth.com/chatrwd. However, most of the RAG studies in medicine use the conventional architecture with only one round of retrieval. There have been several attempts to use iterative data refinement for LLM training[30] or RAG[31, 32, 33] in the general domain. Nevertheless, similar ideas have not yet been explored in the medical domain. To the best of our knowledge, our study presents the first approach and evaluations on incorporating follow-up queries in RAG for medicine.
@@ 4.2 Main results
Table 4.1 shows the comparison results of i-MedRAG and other baseline approaches on MedQA using GPT-3.5. Official scores reported in previous research are used for a fair comparison. While methods with few-shot learning or model fine-tuning tend to perform better than the ones in a zero-shot setting, our i-MedRAG set a state-of-the-art performance of GPT-3.5 on MedQA without any training samples or parameter tuning. Among zero-shot approaches, i-MedRAG (69.68%) has a significant performance improvement ($p<0.05$) compared to the previous best record achieved by MedRAG (66.61%).
The results of generalizing i-MedRAG to more LLMs and data are presented in Table 4.2. We compare i-MedRAG with our implemented CoT and MedRAG to see if i-MedRAG can bring a consistent improvement of LLM performance in medical question answering. For all experiments with i-MedRAG, we tune the hyperparameters on a validation set of 100 samples and then report its scores on the test set. Similar to the results on GPT-3.5, the open-source Llama-3.1-8B also shows improved performance on MedQA with the help of i-MedRAG. While Llama-3.1-8B shows a close performance to GPT-3.5 in CoT and MedRAG settings, its performance is significantly improved with i-MedRAG, achieving an accuracy of 75.02%. The improved performance of GPT-3.5 and Llama-3.1-8B on MMLU-Med also demonstrates the generalizability of i-MedRAG to more medical data. As medical questions in MMLU-Med are less complex than the USMLE questions in MedQA, follow-up queries may not be necessary to find relevant information for the given question. Thus, it can be observed that the improvement by i-MedRAG compared to MedRAG is less significant in MMLU-Med than in MedQA.
@@
Performance of i-MedRAG on different LLMs and datasets. “Acc.” denotes the accuracy. “$\Delta$” shows the relative performance improvement compared with CoT. Model Method MedQA-USMLE MMLU-Med Average Acc. $\Delta$ Acc. $\Delta$ Acc. $\Delta$ GPT-3.5-Turbo CoT 65.04 +0.00% 72.91 +0.00% 68.98 +0.00% GPT-3.5-Turbo MedRAG 66.61 +2.41% 75.48 +3.52% 71.05 +3.00% GPT-3.5-Turbo i-MedRAG 69.68 +7.13% 75.85 +4.03% 72.77 +5.49% Llama-3.1-8B CoT 64.73 +0.00% 77.23 +0.00% 70.98 +0.00% Llama-3.1-8B MedRAG 66.54 +2.80% 78.05 +1.06% 72.30 +1.86% Llama-3.1-8B i-MedRAG 73.61 +13.72% 78.42 +1.54% 76.02 +7.10%
@@ 4.3 Scaling with iterations and queries
Figure 3: i-MedRAG performance on MedQA and MMLU-Med with different hyperparameters.
@@ 4.4 Case studies
Table 4.4 shows the predictions of GPT-3.5 on a MedQA question with different prompt engineering approaches. The question asks about the mechanism of the drug for transitional cell carcinoma of the bladder which causes hearing loss. To solve the problem, it is important to find the exact drug and then figure out how it causes the mentioned symptoms. However, the CoT result shows that GPT-3.5 does not inherently contain sufficient medical knowledge to solve this problem. Instead of inferring the described drug, GPT-3.5 with CoT directly hallucinates a wrong option as the answer. While free radicals are relevant to hearing loss, their connection to the disease of the patient is unclear and not discussed. Compared to CoT which solely relies on the internal knowledge of LLMs, MedRAG provides an opportunity for LLMs to augment their answer generation with external medical knowledge. Nevertheless, the model output shows that the MedRAG system fails to retrieve useful information about the drug from medical corpora. Given the complex problem description, it is difficult for text retrievers to find the asked mechanism without knowing the drug.
@@ 5.1 Limitations
Another limitation is the selection of hyperparameter values for optimal performance. As shown in Figure 3, different LLMs can have different hyperparameter settings for their optimal performance. Even for the same LLM, its optimal hyperparameters can vary based on the medical questions being processed. Thus, it is non-trivial to find the optimal hyperparameters of i-MedRAG for a new medical task, which may be inefficient for real-world deployments.
@@ 5.2 Future work
Given the limitations of i-MedRAG, we consider several potential future directions that could further improve the performance of retrieval-augmented generation for medicine. The first direction is the automation of hyperparameter selection in i-MedRAG. To reduce the laborious process of hyperparameter selection, one may use an LLM agent to dynamically determine how many follow-up queries should be asked each iteration. This can improve the efficiency and flexibility of the hyperparameter selection process. Another future direction is to improve the performance of i-MedRAG with few-shot demonstrations. While few-shot CoT prompting is demonstrated to perform better than the zero-shot counterpart[43], it is not easy to adapt such strategies to i-MedRAG as the reasoning process can be dynamically affected by the use of external corpora and retrievers. Investigating how i-MedRAG can benefit from one or few-shot samples could be a potential direction to further enhance its performance on medical question answering. More quantitative analysis can also be performed to examine the error types of i-MedRAG compared to existing methods.